. Find the electric field a distance z for an infinitely long straight wire, which carries a uniform line charge λ 


1
Expert's answer
2021-07-05T17:46:16-0400

We know that

Electric flux"\phi=E.A"

"\int E.A=\frac{q}{\epsilon_0}"

"E{2\pi rx}=\frac{q}{\epsilon_0}"

"\lambda=\frac{q}{x}"

"q=\lambda x"

"E=\frac{q}{2\pi r \epsilon _0x}"

"E=\frac{\lambda x}{2\pi rx\epsilon_0}"

"E=\frac{\lambda}{2 \pi r\epsilon_0}"